OpenStudy (anonymous):

Write as a decimal: 203%

OpenStudy (mathstudent55):

% = 1/100 = 0.01 203 = 203 * 1/100 = 203 * 0.01

OpenStudy (anonymous):

2.03

OpenStudy (anonymous):

2.03